An investor invested a total amount of $3100.00 in two mutual funds. One mutual fund earned 4% while the other mutual fund earned 2%. If the investors profit was $98.00 how much was invested in each fund?

Answer by mananth(16946)

Part I 2.00% per annum ------------- Amount invested =x
Part II 4.00% per annum ------------ Amount invested = y =
3100
Interest----- 98

Part I 2.00% per annum ---x
Part II 4.00% per annum ---y
Total investment
x + 1 y= 3100 -------------1
Interest on both investments
2.00% x + 4.00% y= 98
Multiply by 100
2 x + 4 y= 9800.00 --------2
Multiply (1) by -2
we get
-2 x -2 y= -6200.00
Add this to (2)
0 x 2 y= 3600
divide by 2
y = 1800
Part I 2.00% $ 1300
Part II 4.00% $ 1800

CHECK
1300 --------- 2.00% ------- 26.00
1800 ------------- 4.00% ------- 72.00
Total -------------------- 98.00

m.ananth@hotmail.ca